OpenStudy (shortypants):

A radiology specialist uses the radioactive substance​ iodine-131 to diagnose conditions of the thyroid gland. His hospital currently has a 24​-gram supply of​ iodine-131. The following table gives the number of grams remaining after a specified number of days. t​ (Number of Days Starting from A 24​-Gram Supply of​ Iodine-131) 0 1 2 3 4 5 6 ​N, Number of Grams of​ Iodine-131 Remaining from A​ 20-Gram Supply 24.00 22.39 20.89 18.19 16.97 16.97 15.83

OpenStudy (shortypants):

c. Write an exponential decay formula for​ N, the number of grams of​ iodine-131 remaining, in terms of​ t, the number of days from the current supply of 24 grams. The formula is N= ​(Use integers or decimals for any numbers in the​ expression.) d. Determine the number of grams of​ iodine-131 remaining from a 24​-gram supply after 22 months ​(60 ​days). N≈ ​(Type an integer or decimal rounded to the nearest hundredth as​ needed.)

OpenStudy (shortypants):

How long will it take for​ iodine-131 to decrease to half its original​ value? It will take about _______days for​ iodine-131 to to decrease to half its original value. ​(Round to the nearest integer as​ needed) Explain how to determine the answer in part f graphically.

OpenStudy (evoker):

General Equation is Amount remaining = Starting Amount * (Decay Rate) ^ t. Solve for the Decay rate by plugging in two points to solve or basically it is the later value divided by the earlier value of two consecutive readings.

OpenStudy (wolf1728):

Here are 4 half-life formulas and here is a link to half-life calculations: http://www.1728.org/halflife.htm

OpenStudy (wolf1728):

Use this equation: remaining amount or ending amount = beginning amount / 2^n where n= number of half-lives n = elapsed time/half-life Note: the half-life of Iodine 131 is 8.0197 days https://en.wikipedia.org/wiki/Iodine-131 Let's calculate after 1 day remaining amount = 24/ 2^n where n = 1/8.0197 remaining amount = 24/2^(1/8.0197) remaining amount = 24/2^0.1246929436 remaining amount = 24/1.0902756589 remaining amount after 1 day= 22.0127816345 grams